# MIRAGE: A Benchmark for Multimodal Information-Seeking and Reasoning in Agricultural Expert-Guided Conversations

### v1

@@ E.1.2 Standard MG Benchmark Results
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Scores represent performance on four key metrics (Accuracy / Relevance / Completeness / Parsimony) on a 0-4 scale. Bold purple values highlight the best performance on each metric within each benchmark.
@@ E.3 With/Without Meta Data Results
Tables 9 and 10 contrast model performance when the question is presented with versus without the user’s location and timestamp. For the Identification benchmark (Table 9), adding metadata yields only modest absolute gains ( pp ID%) and virtually no change in reasoning quality (). The strongest closed-source model (GPT-4.1) and the largest open-source Qwen-72B both extract a small benefit (+1.6 and +0.6 pp, respectively), whereas lighter counterparts (e.g., GPT-4.1-mini, Gemma-3B) see negligible or even negative shifts, suggesting that metadata can introduce noise unless the model possesses sufficient capacity to interpret it.
@@ E.4 Fine-Turning Results
Effect of LoRA fine-tuning.
@@
Figure 10 tracks identification accuracy and reasoning accuracy on seen vs. unseen entities, as Qwen2.5-VL models undergo progressively longer LoRA fine-tuning. For both the 32 B (left) and 7 B (right) variants, the bulk of the improvement is realised within the first four epochs: ID accuracy on seen entities rises from 32.9% to 37.6% for 32 B and from 27.7% to 34.8% for 7 B. Beyond epoch 4 the gains plateau or slightly regress, hinting at diminishing returns and possible over-fitting to the fine-tuning set. Reasoning accuracy follow a similar but more muted trend, increasing by at most 0.2–0.3 points before flattening.
The persistently low curves for unseen entities are unsurprising. Identification requires the model to emit an explicit entity name; if that name never appeared in the fine-tuning set, fine-tuning does not help.
E.5 Category-Wise Breakdown Results
[…]
E.5.1 Plant Identification Reults (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
@@ E.5.1 Plant Identification Reults (MMST Standard)
| gpt-4.1 | 48.7 | 3.13 | 49.0 | 2.89 | 45.9 | 3.28 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | ||||||
| Qwen2.5-VL-32B-Instruct | ||||||
| Qwen2.5-VL-72B-Instruct | ||||||
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | ||||||
| gemma-3-27b-it | ||||||
| InternVL3-2B | ||||||
| InternVL3-8B | ||||||
| InternVL3-14B | ||||||
| InternVL3-38B | ||||||
| InternVL3-78B | ||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
E.5.2 Insect and Pest Identification Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
@@ E.5.2 Insect and Pest Identification Results (MMST Standard)
| gpt-4.1 | 34.9 | 2.93 | 34.9 | 2.60 | 32.6 | 3.01 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | ||||||
| Qwen2.5-VL-32B-Instruct | ||||||
| Qwen2.5-VL-72B-Instruct | ||||||
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | ||||||
| gemma-3-27b-it | ||||||
| InternVL3-2B | ||||||
| InternVL3-8B | ||||||
| InternVL3-14B | ||||||
| InternVL3-38B | ||||||
| InternVL3-78B | ||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
E.5.3 Plant Disease Identification Results (MMST Standard)
| Model | DeepSeek-R1-Distill | Qwen3-32B | Phi-4-reasoning | |||
@@ E.5.3 Plant Disease Identification Results (MMST Standard)
| gpt-4.1 | 45.7 | 3.06 | 44.6 | 2.62 | 46.0 | 2.97 |
| gpt-4.1-mini | ||||||
| gpt-4o | ||||||
| gpt-4o-mini | ||||||
| claude-3-7-sonnet | ||||||
| claude-3-5-sonnet | ||||||
| claude-3-haiku | ||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||
| llava-v1.6-mistral-7b-hf | ||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||
| Qwen2.5-VL-3B-Instruct | ||||||
| Qwen2.5-VL-7B-Instruct | ||||||
| Qwen2.5-VL-32B-Instruct | ||||||
| Qwen2.5-VL-72B-Instruct | ||||||
| gemma-3-4b-it | ||||||
| gemma-3-12b-it | ||||||
| gemma-3-27b-it | ||||||
| InternVL3-2B | ||||||
| InternVL3-8B | ||||||
| InternVL3-14B | ||||||
| InternVL3-38B | ||||||
| InternVL3-78B | ||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Bold purple values denote the best performance in each column.
E.5.4 Plant Disease Management Results (MMST Standard)
@@
| gpt-4.1 | 3.05 | 3.54 | 3.31 | 2.87 | 3.17 | 3.54 | 2.90 | 2.92 | 3.15 | 3.53 | 3.12 | 3.05 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source models in red/light red, open-source models in blue/light blue. Scores are given on a 0–4 scale for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple numbers denote the best performance for each metric within a column block.
E.5.5 Insect and Pest Management Results (MMST Standard)
@@
| gpt-4.1 | 3.05 | 3.51 | 3.33 | 2.87 | 3.18 | 3.53 | 2.98 | 2.98 | 3.11 | 3.52 | 3.15 | 3.06 |
| gpt-4.1-mini | ||||||||||||
| gpt-4o | ||||||||||||
| gpt-4o-mini | ||||||||||||
| nova-pro | ||||||||||||
| claude-3-7-sonnet | ||||||||||||
| claude-3-5-sonnet | ||||||||||||
| claude-3-haiku | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | ||||||||||||
| llava-v1.6-mistral-7b-hf | ||||||||||||
| llava-onevision-qwen2-7b-ov-hf | ||||||||||||
| Qwen2.5-VL-3B-Instruct | ||||||||||||
| Qwen2.5-VL-7B-Instruct | ||||||||||||
| Qwen2.5-VL-32B-Instruct | ||||||||||||
| Qwen2.5-VL-72B-Instruct | ||||||||||||
| gemma-3-4b-it | ||||||||||||
| gemma-3-12b-it | ||||||||||||
| gemma-3-27b-it | ||||||||||||
| InternVL3-2B | ||||||||||||
| InternVL3-8B | ||||||||||||
| InternVL3-14B | ||||||||||||
| InternVL3-38B | ||||||||||||
| InternVL3-78B | ||||||||||||
Models are color-coded by type: closed-source (red) and open-source (blue). Scores range from 0-4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple indicates the best score in each column block.
E.5.6 Plant Care and Gardening Guidance Results (MMST Standard)

### v2

@@ E.1.2 Standard MG Benchmark Results
Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Scores represent performance on four key metrics (Accuracy / Relevance / Completeness / Parsimony) on a 0‑4 scale. Bold purple values highlight the best performance on each metric within each benchmark.
@@ E.3 With/Without Meta Data Results
Tables 9 and 10 compare model performance when each question is presented with versus without the user’s location and timestamp. Metadata is incorporated by appending this spatiotemporal context in natural-language form to the image–text input during inference. For the Identification benchmark, gains are modest (,pp ID%) with negligible changes in reasoning quality (). Stronger models (GPT-4.1, Qwen-72B) show slight improvements (+1.6 and +0.6,pp), whereas smaller ones (GPT-4.1-mini, Gemma-3B) show minimal or negative shifts, indicating that metadata can add noise without sufficient model capacity.
@@ E.4 Fine-Turning Results
Effect of LoRA fine‐tuning.
@@
Figure 10 tracks identification accuracy and reasoning accuracy on seen vs. unseen entities, as Qwen2.5-VL models undergo progressively longer LoRA fine-tuning. For both the 32 B (left) and 7 B (right) variants, the bulk of the improvement is realised within the first four epochs: ID accuracy on seen entities rises from 32.9% to 37.6% for 32 B and from 27.7% to 34.8% for 7 B. Beyond epoch 4 the gains plateau or slightly regress, hinting at diminishing returns and possible over-fitting to the fine-tuning set. Reasoning accuracy follow a similar but more muted trend, increasing by at most 0.2–0.3 points before flattening. The persistently low curves for unseen entities are unsurprising. Identification requires the model to emit an explicit entity name; if that name never appeared in the fine-tuning set, fine-tuning does not help.
Appendix F Error Analysis
We conducted a quantitative error analysis to complement our main evaluation results for deeper insights into model failure modes. We performed a stratified manual error review across three scenarios: (i) GPT-4.1 failures ( of 2,395), (ii) Qwen2.5-VL-32B failures where GPT-4.1 succeeds ( of 1,078), and (iii) joint failures ( of 300). We focused on instances with zero identification accuracy and compared the failure modes and root causes between the proprietary GPT-4.1 and the open-source Qwen2.5-VL-32B. The observed errors cluster into two tiers: Fundamental Domain Challenges and Open-Source Model Systematic Failures.
F.1 Tier 1: Fundamental Domain Challenges
These represent the current boundaries of what even the most advanced LVLMs can achieve in agricultural reasoning and consultation contexts. Failures in this tier occur in GPT-4.1, the best-performing model in our evaluation, indicating inherent limitations of current LVLMs in agricultural expert reasoning. Percentages reflect the relative frequency of each error type within sampled failures
Specialist Knowledge Gaps (35% of cases): These involve missing domain-specific understanding of biological phenomena requiring deep expertise beyond general training data. The LVLMs fail to recognize or reason about biological processes. Example: GPT-4.1 could not interpret “appleleaf blister mite” damage or its developmental process.
Complex Diagnostic Reasoning (10% of cases): These involve correct symptom observation but incorrect causal inference or disease conclusion. Despite accurate visual analysis, the models fail to reason over underlying biological causality. Example: Observing insect damage but misdiagnosing the pest species.
Finegrained visual understanding (55% of cases): Majority of the errors occured due to model’s inability to pick up finegrained visual cues and distinguishing factors among different adjacent biological species or part occlusions of the main entity in images.
F.2 Tier 2: Open-Source Model Systematic Failures
These failure types occur primarily in Qwen2.5-VL-32B but not in GPT-4.1, revealing systematic capability gaps between open-source and proprietary LVLMs.
Vision-Language Integration Issues (40% of cases): Failures in integrating visual cues with domain-specific terminology lead to misidentification among visually similar species. Domain-specific finetuning mitigates this issue. Example: Qwen2.5-VL-32B confusing a longhorn beetle with a June beetle.
Diagnostic Hedging Bias (25% of cases): The open-source model frequently exhibits hedging behavior, avoiding decisive expert-level conclusions.
Generalist Reasoning Bias (20% of cases): The model applies broad, non-specific reasoning, listing multiple generic possibilities instead of providing a targeted diagnosis.
F.3 Quantitative Error Pattern Summary
| Failure Category | Percentage | Model |
| Tier 1: Fundamental Domain Challenges (GPT-4.1 Failures) | ||
| Specialist Knowledge Gaps | 35% | GPT-4.1 |
| Complex Diagnostic Reasoning | 10% | GPT-4.1 |
| Finegrained visual understanding | 55% | GPT-4.1 |
| Tier 2: Open-Source Model Systematic Failures (Qwen2.5-VL-32B vs. GPT-4.1) | ||
| Vision-Language Integration Issues | 40% | Qwen2.5-VL-32B |
| Diagnostic Hedging Bias | 25% | Qwen2.5-VL-32B |
| Generalist Reasoning Bias | 20% | Qwen2.5-VL-32B |
Critical Finding: In joint failures, GPT-4.1 still demonstrates superior reasoning quality in 60% of cases, reflecting deeper integration of visual and textual cues and more expert-like reasoning patterns, even when incorrect in final identification.
F.4 Performance on Common and Rare Entities
The MIRAGE dataset exhibits a long-tail entity distribution, where models consistently perform worse on rare entities than on common ones. This imbalance exposes a central challenge for agricultural AI: models must recognize thousands of rare species that occur infrequently in training data. Addressing this limitation will require domain adaptation and long-tail learning approaches beyond standard large-scale training.
| Model | DeepSeek-R1 | Qwen3-32B | Phi-4-reasoning |
| GPT-4.1 | 53.1, 3.15 | 52.7, 2.84 | 52.2, 3.20 |
| GPT-4.1-mini | 42.1, 2.82 | 41.3, 2.53 | 40.5, 2.95 |
| Qwen2.5-VL-32B-Instruct | 32.9, 2.62 | 32.2, 2.23 | 31.6, 2.62 |
| Qwen2.5-VL-72B-Instruct | 38.1, 2.65 | 38.4, 2.27 | 37.1, 2.63 |
| Model | DeepSeek-R1 | Qwen3-32B | Phi-4-reasoning |
| GPT-4.1 | 38.5, 3.01 | 38.9, 2.73 | 35.4, 3.14 |
| GPT-4.1-mini | 32.1, 2.76 | 30.4, 2.50 | 26.8, 2.94 |
| Qwen2.5-VL-32B-Instruct | 21.1, 2.47 | 20.4, 2.14 | 18.2, 2.53 |
| Qwen2.5-VL-72B-Instruct | 25.5, 2.56 | 24.5, 2.18 | 22.1, 2.57 |
Building on these findings, we outline a roadmap for advancing multimodal large vision-language models (VLMs) toward robust, interactive agricultural consultation systems. Future versions MIRAGE should integrate agentic capabilities that enable models to proactively utilize time- and location-based context—for example, adapting responses based on seasonal patterns, regional crop profiles, or current environmental conditions. This contextual grounding will allow evaluations to move beyond static reasoning and toward dynamic, situation-aware dialogue.
To achieve this, future VLM development should emphasize the following directions:
Contextual and Temporal Grounding: Incorporate temporal signals (e.g., crop cycles, weather timelines) and geospatial information (e.g., soil types, local pest occurrences) to support temporally and regionally coherent reasoning in consultation tasks.
Interactive Conversational Reasoning: Extend MIRAGE’s multi-turn benchmark to simulate fully interactive dialogues, where models must perform clarification, conversational repair, and feedback incorporation, key aspects of real-world agricultural advisory systems.
Domain-Specific Knowledge Integration: Bridge VLMs with structured agronomic knowledge graphs and expert-curated datasets to reduce specialist knowledge gaps identified in our error analysis.
Adaptive Visual-Linguistic Understanding: Improve VLMs’ capacity to detect fine-grained visual cues and align them with domain terminology, addressing the vision-language integration failures seen in open-source models.
Long-Tail Entity Recognition: Develop targeted fine-tuning and retrieval-augmented methods for handling rare species and diseases that dominate real-world agricultural problem distributions.
These extensions will move MIRAGE and future VLMs from static perception and reasoning benchmarks toward fully interactive, contextually grounded expert systems. Such systems could ultimately assist human agricultural specialists in diagnosing issues, providing adaptive recommendations, and integrating multimodal data streams for decision support in diverse agricultural environments.
F.5 Category-Wise Breakdown Results
F.5.1 Plant Identification Reults (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||
@@ F.5.1 Plant Identification Reults (MMST Standard)
| gpt‑4.1 | 48.7 | 3.13 | 49.0 | 2.89 | 45.9 | 3.28 |
| gpt‑4.1‑mini | ||||||
| gpt‑4o | ||||||
| gpt‑4o‑mini | ||||||
| claude‑3‑7‑sonnet | ||||||
| claude‑3‑5‑sonnet | ||||||
| claude‑3‑haiku | ||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||
| gemma‑3‑4b‑it | ||||||
| gemma‑3‑12b‑it | ||||||
| gemma‑3‑27b‑it | ||||||
| InternVL3‑2B | ||||||
| InternVL3‑8B | ||||||
| InternVL3‑14B | ||||||
| InternVL3‑38B | ||||||
| InternVL3‑78B | ||||||
Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Bold purple values denote the best performance in each column.
F.5.2 Insect and Pest Identification Results (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||
@@ F.5.2 Insect and Pest Identification Results (MMST Standard)
| gpt‑4.1 | 34.9 | 2.93 | 34.9 | 2.60 | 32.6 | 3.01 |
| gpt‑4.1‑mini | ||||||
| gpt‑4o | ||||||
| gpt‑4o‑mini | ||||||
| claude‑3‑7‑sonnet | ||||||
| claude‑3‑5‑sonnet | ||||||
| claude‑3‑haiku | ||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||
| gemma‑3‑4b‑it | ||||||
| gemma‑3‑12b‑it | ||||||
| gemma‑3‑27b‑it | ||||||
| InternVL3‑2B | ||||||
| InternVL3‑8B | ||||||
| InternVL3‑14B | ||||||
| InternVL3‑38B | ||||||
| InternVL3‑78B | ||||||
Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Bold purple values denote the best performance in each column.
F.5.3 Plant Disease Identification Results (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||
@@ F.5.3 Plant Disease Identification Results (MMST Standard)
| gpt‑4.1 | 45.7 | 3.06 | 44.6 | 2.62 | 46.0 | 2.97 |
| gpt‑4.1‑mini | ||||||
| gpt‑4o | ||||||
| gpt‑4o‑mini | ||||||
| claude‑3‑7‑sonnet | ||||||
| claude‑3‑5‑sonnet | ||||||
| claude‑3‑haiku | ||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||
| gemma‑3‑4b‑it | ||||||
| gemma‑3‑12b‑it | ||||||
| gemma‑3‑27b‑it | ||||||
| InternVL3‑2B | ||||||
| InternVL3‑8B | ||||||
| InternVL3‑14B | ||||||
| InternVL3‑38B | ||||||
| InternVL3‑78B | ||||||
Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Bold purple values denote the best performance in each column.
F.5.4 Plant Disease Management Results (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.4 Plant Disease Management Results (MMST Standard)
| gpt‑4.1 | 3.05 | 3.54 | 3.31 | 2.87 | 3.17 | 3.54 | 2.90 | 2.92 | 3.15 | 3.53 | 3.12 | 3.05 |
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | ||||||||||||
| gpt‑4o‑mini | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source models in red/light red, open‑source models in blue/light blue. Scores are given on a 0–4 scale for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple numbers denote the best performance for each metric within a column block.
F.5.5 Insect and Pest Management Results (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||
@@ F.5.5 Insect and Pest Management Results (MMST Standard)
| gpt‑4.1 | 3.05 | 3.51 | 3.33 | 2.87 | 3.18 | 3.53 | 2.98 | 2.98 | 3.11 | 3.52 | 3.15 | 3.06 |
| gpt‑4.1‑mini | ||||||||||||
| gpt‑4o | ||||||||||||
| gpt‑4o‑mini | ||||||||||||
| nova‑pro | ||||||||||||
| claude‑3‑7‑sonnet | ||||||||||||
| claude‑3‑5‑sonnet | ||||||||||||
| claude‑3‑haiku | ||||||||||||
| Llama‑4‑Scout‑17B‑16E‑Instruct | ||||||||||||
| llava‑v1.6‑mistral‑7b‑hf | ||||||||||||
| llava‑onevision‑qwen2‑7b‑ov‑hf | ||||||||||||
| Qwen2.5‑VL‑3B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑7B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑32B‑Instruct | ||||||||||||
| Qwen2.5‑VL‑72B‑Instruct | ||||||||||||
| gemma‑3‑4b‑it | ||||||||||||
| gemma‑3‑12b‑it | ||||||||||||
| gemma‑3‑27b‑it | ||||||||||||
| InternVL3‑2B | ||||||||||||
| InternVL3‑8B | ||||||||||||
| InternVL3‑14B | ||||||||||||
| InternVL3‑38B | ||||||||||||
| InternVL3‑78B | ||||||||||||
Models are color‑coded by type: closed‑source (red) and open‑source (blue). Scores range from 0‑4 for Accuracy (Acc), Relevance (Rel), Completeness (Comp), and Parsimony (Pars). Bold purple indicates the best score in each column block.
F.5.6 Plant Care and Gardening Guidance Results (MMST Standard)
| Model | DeepSeek‑R1‑Distill | Qwen3‑32B | Phi‑4‑reasoning | |||||||||